Liz commutes 30 mi. to her job each day. She finds that if she drives 10 mi/h faster it takes her 6 minutes less to get to work. Find her new speed.
Answer by mananth(16946)

1/2 hour to work place every day
speed = 30 mph
...
if she increases her speed by 10 mph her speed will be 30+10 = 40 mph.
...
let distance to work place be x
..
time taken with present speed - time taken with increased speed = 1/10 hours
..
t= d/r
..
x/30 - x/(40)= 1/6
LCD = 120
(4x-3x)/120 = 1/6
x/120 =1/6
6x=120
/6
x= 20 miles distance to work
..
m.ananth@hotmail.ca